# Convergence Rates for Stochastic Proximal and Projection Estimators

### v2

@@ 1. Introduction
Proximal mappings and metric projections are fundamental primitives in variational analysis and optimization. Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$. When $f=\iota_{C}$ is the indicator of a closed convex set $C$, the proximal mapping reduces to the metric projection $\operatorname{proj}_{C}(x)$. In many applications, however, evaluating $\operatorname{prox}_{\lambda f}(x)$ or $\operatorname{proj}_{C}(x)$ exactly is computationally expensive, and in black-box settings only function values (possibly noisy) are available.
This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting. For $\delta>0$, define the barycenter
@@
The estimators (1) arise naturally from Laplace-type asymptotics for integrals with kernel $\exp(-g/\delta)$, where $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$. Two recent strands of work developed this idea.
First, motivated by Hamilton-Jacobi equations, [9] proposed the HJ-Prox method and proved that, for $\rho$-weakly convex objectives under mild regularity, the barycenter converges to the proximal point as $\delta\to 0^{+}$. Second, [10] places the relationship between infimal convolutions and self-normalized Laplace approximation front and center, extending the construction beyond the quadratic kernel and establishing asymptotic validity under weak local assumptions; in particular, their framework includes smoothed set projections $p_{\delta}(x)=\mathbb{E}[Y\mid Y\in C]$ and proves $p_{\delta(x)}\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$ under mild local geometric conditions near the projection point.
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
@@ 2. Mathematical Preliminaries
Throughout this paper, we work in the finite-dimensional Hilbert space $\mathbb{R}^{n}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ and the associated norm $\|\cdot\|$. The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$. We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
For a ser $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
@@
Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$. Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{\mathbb{R}}\cup\{+\infty\}$, the domain of $g$ is
@@
The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to{\mathbb{R}}_{\infty}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
@@
the expectation of a given function $w$ with respect to this probability measure.
The following result is known as Fradelizi’s inequality (see [6])
Lemma 2.1.
Let $h:\mathbb{R}^{n}\to[0,\infty)$ be a log-concave function, not identically zero, such that $\int_{\mathbb{R}^{n}}h(y)\,dy<\infty$. Define its barycenter by
$b_{h}:=\frac{\int_{\mathbb{R}^{n}}y\,h(y)\,dy}{\int_{\mathbb{R}^{n}}h(y)\,dy}.$
Then $b_{h}$ is well-defined and
$h(b_{h})\ \geq\ e^{-n}\,\sup_{z\in\mathbb{R}^{n}}h(z).$
3. Nonsmooth results
Let $f\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The aim of this section is to prove that, for any $0<\lambda<1/\rho$,
$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[y\cdot\exp(-f(y)/\delta)]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}[\exp(-f(y)/\delta)]}\to\operatorname{prox}_{\lambda f}(x)\textrm{ as }\delta\to 0^{+}.$
The next result provides a rate of convergence of the barycenter toward the projection.
@@ 3. Nonsmooth results
Claim 1: $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
Proof of Claim 1: Since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption). On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e. and belongs to $W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D))$. Consequently, $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, and for a.e. $y\in\operatorname{int}(D)$,
@@
Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap\mathbb{B}_{R}$. Since
$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }\mathbb{B}_{R}$
and $e^{-g(y)/\delta}\leq e^{-\inf_{\mathbb{B}_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
@@
Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$. For $R>0$, apply the divergence theorem on $\Omega_{R}$:
@@
$\int_{\partial\Omega_{R_{k}}}\langle F,\nu\rangle\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
@@
$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad\langle F(y),\nu_{D}(y)\rangle\geq 0.$
Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{1}$ is bounded in absolutely value by
@@
hence the $\partial\mathbb{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
$0$ $\leq\limsup_{k\to\infty}\int_{\mathbb{B}_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap\mathbb{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial\mathbb{B}_{R_{k}}\cap D}F\cdot\nu_{\mathbb{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\,\nu\,d\mathcal{H}^{n-1}$
@@
Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the claim.∎
Next, by strong convexity and almost-everywhere differentiability, we obtain
@@
hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen inequality,
@@
The following example shows that the bound obtained in Theorem 3.1 is sharp.
@@
Fix $n\geq 2$, take $\lambda>0$, set $\mu=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\}.$
Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]m_{\alpha,n},$
@@
$m_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$. In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $m_{\alpha,n}\in[\cos\alpha,1]$ and hence $m_{\alpha,n}\to 1$ as $\alpha\searrow 0$. Furthermore,
$\mathbb{E}[\chi_{n}]\sqrt{n}\to 1\textrm{ as }n\to\infty.$
@@
$\|p_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
@@ 4. Smooth results
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\,\delta.$
@@
Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and
From this inequality we get that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
Given a function $w(\cdot)$, we define
@@
Proof of Claim 1: We observe that
@@
$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{3nL}{2\mu^{2}}\delta.$
@@
where we have used that $H\succcurlyeq\mu I$ and $\|H^{-1}\|\leq 1/\mu$. Moreover,
@@
(6) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\quad\Rightarrow\quad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu}.$
Next, again since $h_{\delta}=e^{-g/\delta}$ is log-concave, applying the Lemma 2.1 to $h_{\delta}$ yields
(7) $g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\leq n\delta.$
By $\mu$-strong convexity of $g$,
$g(m_{\delta}(x))-g(\operatorname{prox}_{\lambda f}(x))\geq\frac{\mu}{2}\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2},$
and combining with (7) gives
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{2n\delta}{\mu}.$
@@
$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{2n\delta}{\mu}=\frac{3n\delta}{\mu}.$
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{3n\delta}{\mu}=\frac{3nL}{2\mu^{2}}\delta,$

### v3

@@ 1. Introduction
Proximal mappings and metric projections are fundamental concepts in variational analysis and optimization. Introduced by J.-J. Moreau in [9], proximal mappings have since become central tools in modern optimization, notably through proximal algorithms and operator-splitting methods and their many variants (see, e.g., [2]). Given a proper function $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and $\lambda>0$, the proximal point $\operatorname{prox}_{\lambda f}(x)$ is the unique minimizer of $y\mapsto f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$ whenever $f$ is $\rho$-weakly convex and $\lambda\in(0,1/\rho)$. When $f=\iota_{C}$ is the indicator of a closed convex set $C$, the proximal mapping reduces to the metric projection $\operatorname{proj}_{C}(x)$. In many applications, however, evaluating $\operatorname{prox}_{\lambda f}(x)$ or $\operatorname{proj}_{C}(x)$ exactly is computationally expensive, and in black-box settings only function values (possibly noisy) are available.
This paper studies a family of stochastic (zeroth-order) estimators for proximal points and projections based on Gaussian perturbations and exponential reweighting. For $\delta>0$ and $\lambda>0$, we define the barycenter
@@
The estimators (1) arise naturally from Laplace-type asymptotics for integrals with kernel $\exp(-g/\delta)$, where $g(y):=f(y)+\frac{1}{2\lambda}\|x-y\|^{2}$. Two recent strands of work developed this idea. First, motivated by Hamilton-Jacobi equations, [10] proposed the HJ-Prox method and proved that, for $\rho$-weakly convex objectives under mild regularity, the barycenter converges to the proximal point as $\delta\to 0^{+}$. Second, [11] places the relationship between infimal convolutions and self-normalized Laplace approximation front and center, extending the construction beyond the quadratic kernel and establishing asymptotic validity under weak local assumptions; in particular, their framework includes smoothed set projections $p_{\delta}(x)=\mathbb{E}[Y\mid Y\in C]$ and proves $p_{\delta(x)}\to\operatorname{proj}_{C}(x)$ as $\delta\to 0^{+}$ under mild local geometric conditions near the projection point.
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta,\qquad\mu=\frac{1}{\lambda}-\rho,$
@@ 2. Mathematical Preliminaries
Throughout this paper, we work in the finite-dimensional Hilbert space $\mathbb{R}^{n}$ endowed with the inner product $\langle\cdot,\cdot\rangle$ and the associated norm $\|\cdot\|$. The closed unit ball in $\mathbb{R}^{n}$ is denoted by $\mathbb{B}$. The open ball of radius $R$ centered at $x$ is denoted by $B_{R}(x)$. For simplicity, the open ball of radius $R$ centered at 0 is denoted by $B_{R}$. Moreover, $\mathbb{S}^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$.
We use the convention $\exp(+\infty)=+\infty$ and $\exp(-\infty)=0$.
For a set $A\subset\mathbb{R}^{n}$, the indicator function of $A$ is defined by
@@
Moreover, if $A$ is nonempty, closed and convex, the projection of $x\in\mathbb{R}^{n}$ onto $A$ is denoted by $\operatorname{proj}_{A}(x)$. Given a function $g\colon\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}$, the domain of $g$ is
@@
The Moreau envelope of index $\lambda>0$ of a function $g\colon\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is denoted by $\operatorname{e}_{\lambda}g\colon\mathbb{R}^{n}\to\mathbb{R}$ and is defined by
@@
the expectation of a given function $w$ with respect to the probability measure $\sigma_{\delta}$, which is formally defined for every Lebesgue-measurable set $A$ as
$\sigma_{\delta}(A):=\frac{\int_{A}e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
3. Rates of Convergence for Nonsmooth Functions
Let $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ be a $\rho$-weakly convex function with $\rho\geq 0$. The goal of this section is to establish an explicit convergence rate for the map
$m_{\delta}(x):=\frac{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[y\,\exp(-f(y)/\delta)\bigr]}{\mathbb{E}_{y\sim\mathcal{N}(x,\delta\lambda I)}\!\bigl[\exp(-f(y)/\delta)\bigr]},$
and to quantify its approximation of the proximal point:
$m_{\delta}(x)\;\longrightarrow\;\operatorname{prox}_{\lambda f}(x)\qquad\text{as }\delta\downarrow 0.$
It is straightforward to verify that if $f$ is $\rho$-weakly convex for some $\rho\geq 0$ and $\operatorname{dom}f$ has nonempty interior, then $m_{\delta}$ is well-defined for every $\delta>0$ and $0<\lambda<1/\rho$.
The following result provides a convergence rate for this barycenter toward the proximal point.
@@ 3. Rates of Convergence for Nonsmooth Functions
Now, let us show that $\mathbb{E}_{\sigma_{\delta}}\left[\langle y-\operatorname{prox}_{\lambda f}(x),\nabla g(y)\rangle\right]\leq n\delta$.
Indeed, since $y\mapsto g(y)$ is proper, lsc, and convex, its domain is convex; in particular, $D:=\operatorname{dom}f=\operatorname{dom}g$ is convex with nonempty interior (by assumption). On $\operatorname{int}(D)$, every proper lsc convex function is locally Lipschitz, hence differentiable a.e. and belongs to $W^{1,1}_{\operatorname{loc}}(\operatorname{int}(D))$. Consequently, $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, and for a.e. $y\in\operatorname{int}(D)$,
@@
Fix $R>0$ and set $\Omega_{R}:=\operatorname{int}(D)\cap B_{R}$. Since
$\|y-\operatorname{prox}_{\lambda f}(x)\|\leq R+\|\operatorname{prox}_{\lambda f}(x)\|=:C_{R}\textrm{ on }B_{R}$
and $e^{-g(y)/\delta}\leq e^{-\inf_{B_{R}}g/\delta}=:M_{R,\delta}<\infty$, it follows that for a.e. $y\in\Omega_{R}$,
@@
Since $\nabla g\in L^{1}_{\operatorname{loc}}(\operatorname{int}(D);\mathbb{R}^{n})$, we have $\nabla g\in L^{1}(\Omega_{R};\mathbb{R}^{n})$, and then $\operatorname{div}F\in L^{1}(\Omega_{R})$. For $R>0$, apply the divergence theorem to $\Omega_{R}$:
@@
$\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}=\underbrace{\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}}_{=:I_{1}}+\underbrace{\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}}_{=:I_{2}}.$
@@
$\langle y-\operatorname{prox}_{\lambda f}(x),\nu_{D}(y)\rangle\geq 0\quad\Rightarrow\quad F(y)\cdot\nu_{D}(y)\geq 0.$
Hence the $\partial D$-term $I_{1}$ is nonnegative. On the other hand, using (2), the term $I_{2}$ is bounded in absolute value by
@@
hence the $\partial{B}_{R_{k}}$-term $I_{2}$ vanishes as $k\to\infty$. Therefore,
$0$ $\leq\limsup_{k\to\infty}\int_{B_{R_{k}}\cap\partial D}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}$
$=\limsup_{k\to\infty}\Big(\int_{\partial D\cap{B}_{R_{k}}}F\cdot\nu_{D}\,d\mathcal{H}^{n-1}+\int_{\partial{B}_{R_{k}}\cap D}F\cdot\nu_{{B}_{R_{k}}}\,d\mathcal{H}^{n-1}\Big)$
$=\limsup_{k\to\infty}\int_{\partial\Omega_{R_{k}}}F\cdot\nu\,d\mathcal{H}^{n-1}$
@@
Finally, dividing by $\int_{D}e^{-g(y)/\delta}dy$ and rearranging, we obtain the estimate. Next, by strong convexity and almost-everywhere differentiability, we obtain
@@
hence $\mathbb{E}_{\sigma_{\delta}}\left[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}\right]\leq n\delta/\mu$. Finally, by Jensen’s inequality,
@@
The following example shows that the bound obtained in Theorem 3.1 is optimal.
@@
Fix $n\geq 2$, take $\lambda>0$, set $\mu:=1/\lambda$, and let $x=0$. For $\alpha\in(0,\pi/2)$, define the circular cone
$K_{\alpha}:=\{y\in\mathbb{R}^{n}\colon\langle y,e_{1}\rangle\geq\|y\|\,\cos\alpha\},$
where $e_{1}$ denotes the first canonical basis vector. Let $f=\iota_{K_{\alpha}}$ be the indicator function of $K_{\alpha}$. Then $\operatorname{prox}_{\lambda f}(0)=0$. Moreover,
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|=\|m_{\delta}(0)\|=\sqrt{\frac{\delta}{\mu}}\mathbb{E}[\chi_{n}]\mathfrak{m}_{\alpha,n},$
@@
$\mathfrak{m}_{\alpha,n}:=\mathbb{E}\left[U_{1}\mid U_{1}\geq\cos\alpha\right],$
with $U$ uniformly distributed on $\mathbb{S}^{n-1}$ and $U_{1}=\langle U,e_{1}\rangle$. In particular, since $U_{1}\in[\cos\alpha,1]$ on the conditioning event, we have $\mathfrak{m}_{\alpha,n}\in[\cos\alpha,1]$ and hence $\mathfrak{m}_{\alpha,n}\to 1$ as $\alpha\searrow 0$. Furthermore,
$\frac{\mathbb{E}[\chi_{n}]}{\sqrt{n}}\to 1\textrm{ as }n\to\infty.$
@@
$\|m_{\delta}(0)-\operatorname{prox}_{\lambda f}(0)\|\geq(1-\varepsilon)\sqrt{\frac{n\delta}{\mu}}.$
@@ 4. Smooth results
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\,\delta.$
@@
Since $f$ is $\rho$-weakly convex and $0<\lambda<1/\rho$, it follows that $g$ is $\mu$-strongly convex and that (2) holds. In particular, (2) implies that $e^{-g/\delta}$ is integrable for every $\delta>0$, so $m_{\delta}(x)$ is well-defined.
Given a function $w(\cdot)$, recall that
$\mathbb{E}_{\sigma_{\delta}}[w(y)]=\frac{\int_{\mathbb{R}^{n}}w(y)e^{-g(y)/\delta}dy}{\int_{\mathbb{R}^{n}}e^{-g(y)/\delta}dy}.$
@@
Proof of Claim 1: First, note that the Lipschitz continuity of the Hessian of $f$ implies that $\nabla g$ grows at most quadratically. More precisely, there exist constants $a,b>0$ such that
$\|\nabla g(y)\|\leq a\|y\|^{2}+b\qquad\text{for all }y\in\mathbb{R}^{n}.$
In particular, $\nabla g$ is integrable with respect to the probability measure $\sigma_{\delta}$.
Second, we observe that
@@
$0\leq e^{-g(y)/\delta}\leq e^{-g(\operatorname{prox}_{\lambda f}(x))/\delta}\cdot e^{-\frac{\mu}{2\delta}\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}}\textrm{ for all }y\in\mathbb{R}^{n},$
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{nL}{\mu^{2}}\delta.$
@@
where we used that $H\succcurlyeq\mu I$, hence $\|H^{-1}\|_{2}\leq 1/\mu$, where $\|\cdot\|_{2}$ is the induced $\ell_{2}$-operator norm. Moreover, using that $\mathbb{E}_{\sigma_{\delta}}[y]=m_{\delta}(x)$, we get that
@@
(5) $\operatorname{Cov}_{\sigma_{\delta}}(y)\preceq\frac{\delta}{\mu}I\qquad\Rightarrow\qquad\mathbb{E}_{\sigma_{\delta}}[\|y-m_{\delta}(x)\|^{2}]=\operatorname{tr}\left(\operatorname{Cov}_{\sigma_{\delta}}(y)\right)\leq\operatorname{tr}\left(\frac{\delta}{\mu}I\right)=\frac{n\delta}{\mu},$
where $\operatorname{Cov}_{\sigma_{\delta}}$ denotes the covariance matrix of a random vector under the probability measure $\sigma_{\delta}$. Moreover, Theorem 3.1 gives us the estimation
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|^{2}\leq\frac{n\delta}{\mu}.$
@@
$\mathbb{E}_{\sigma_{\delta}}[\|y-\operatorname{prox}_{\lambda f}(x)\|^{2}]\leq\frac{n\delta}{\mu}+\frac{n\delta}{\mu}=\frac{2n\delta}{\mu}.$
@@
$\|m_{\delta}(x)-\operatorname{prox}_{\lambda f}(x)\|\leq\frac{L}{2\mu}\cdot\frac{2n\delta}{\mu}=\frac{nL}{\mu^{2}}\delta,$